# Low-depth quantum symmetrization

### v1

@@ Abstract
Quantum symmetrization is the task of transforming a non-strictly increasing list of $n$ integers into an equal superposition of all permutations of the list (or more generally, performing this operation coherently on a superposition of such lists). This task plays a key role in initial state preparation for first-quantized simulations. A weaker version of symmetrization in which the input list is strictly increasing has been extensively studied due to its application to fermionic systems, but the general symmetrization problem with repetitions in the input list has not been solved previously. We present the first quantum algorithm for the general symmetrization problem. Our algorithm, based on (quantum) sorting networks, uses $O(\log n)$ depth and $O(n\log m)$ ancilla qubits where $m$ is the greatest possible value of the list, enabling efficient simulation of bosonic quantum systems in first quantization. Furthermore, our algorithm can prepare (superpositions of) Dicke states of any Hamming weight in $O(\log n)$ depth using $O(n\log n)$ ancilla qubits. We also propose an $O(\log^{2}n)$-depth quantum algorithm to transform second-quantized states to first-quantized states. Using this algorithm, QFT-based quantum telescope arrays can image brighter photon sources, extending quantum interferometric imaging systems to a new regime.
@@ I Introduction
Symmetry is a central concept in physics. In particular, symmetry plays a key role in quantum simulation Feynman (1982), where a faithful simulation of the many-body dynamics of indistinguishable particles should respect the symmetry of the simulated system. For fermions, due to the Pauli exclusion principle, no two particles can occupy the same mode, and the wave function must be antisymmetrized with respect to permutations of particles, leading to Fermi-Dirac statistics. For bosons, multiple particles can be in the same state, and the wave function must be symmetrized with respect to permutations, leading to Bose-Einstein statistics. A general quantum simulation procedure contains not only Hamiltonian evolution and final output, but also initial state preparation, which is often a challenging step, especially when the initial wave function has some specific structure or symmetry.
@@
In this paper, we study a simple but subtle problem in quantum computing: transforming a list $l$ of non-strictly increasing integers stored in a quantum register into the equal superposition of all permutations of $l$. We call this “symmetrization,” since the output state is symmetric with respect to any permutation of $l$. Symmetrization plays a fundamental role in quantum simulation based on first quantization, an encoding of identical particles as a list of particle locations (whereas second quantization represents the state as a list of occupation numbers of particle locations). Eigenstates in first quantization must be symmetric or antisymmetric if the system is bosonic or fermionic, respectively. Due to significant interest in quantum chemistry (which involves interactions between fermions), a weaker version of (anti)symmetrization where the list is strictly increasing has been extensively studied Abrams and Lloyd (1997); Ward et al. (2009); Berry et al. (2018). In particular, work of Berry et al. (2018) develops a logarithmic-depth algorithm. However, first-quantized simulation of bosons must allow multiple particles to be in the same mode, so the input list may contain repeated integers. Although several previous papers have claimed to handle this more general case Abrams and Lloyd (1997); Ward et al. (2009); Su et al. (2021); Chan et al. (2023); Kosugi et al. (2023); Georges et al. (2024), close examination reveals a subtlety that has not been properly addressed in the literature.
@@
As the main results of this paper, we propose new logarithmic-depth quantum algorithms for symmetrization and a polylogarithmic-depth quantum algorithm for transforming from second to first quantization. We also present three applications of these algorithms. The first application, mentioned above, is initial state preparation in first-quantized simulation of bosons, an unrecognized unsolved problem before our work. The second application is preparing a specific class of symmetric states useful in a variety of quantum information processing tasks, Dicke states, where our algorithm is the first  [^1] We are aware of concurrent work that also gives an algorithm for preparing Dicke states with polylogarithmic depth but uses only polylogarithmically many ancillas Yu et al. (2024). The approaches are very different: while we use sorting networks, the other approach applies a simple sequence of collective rotations and parity measurements. Our approach solves a more general symmetrization problem, while the other approach is simpler, likely performs better in practice, and can perform better when the Hamming weight is lower. to achieve logarithmic total circuit depth using a reasonable number of ancilla qubits. The third application is in optics, where we use our algorithm to improve strategies for long-baseline quantum telescope arrays Gottesman et al. (2012); Khabiboulline et al. (2019a) to process multiple photons at a time, yielding a quantum interferometric imaging system that can take images in more general scenarios than previous designs.
@@
We present background for our results in Section II. More specifically, in Section II.1, we introduce sorting networks and their quantization as the building blocks of symmetrization. Then in Section II.2, we describe the algorithm of Berry et al. (2018), which is used as a subroutine of one of our symmetrization algorithms. In Section II.3, we generalize sorting to different comparison rules and define reversible sorting networks. In Section II.4, we present the quantization of a standard prefix sum computation procedure.
In Section III, we discuss the general symmetrization problem with repetitions. We first explain in Section III.1 why previous methods Abrams and Lloyd (1997); Ward et al. (2009); Berry et al. (2018) are not applicable to the more general symmetrization problem by discussing the subgroup structure of permutations when considering repetitions in the permuted list, and then extend quantum sorting networks to work with repetitions in Section III.2. Using these theoretical tools, we present our new algorithm for symmetrizing a single input list in Section III.3 and then explain how to symmetrize an arbitrary superposition of input lists in Section III.4.
In Section IV, we propose a polylogarithmic-depth converter from a second quantized representation to a sorted list of modes occupied by the particles. This binary encoder, together with the symmetrization algorithm, gives the first polylogarithmic-depth conversion from second quantization to first quantization.
As a topic of independent interest, we demonstrate our design of a quantum interferometric imaging system in Section V, including a pedagogical introduction of quantum telescope arrays that process one photon a time Gottesman et al. (2012); Khabiboulline et al. (2019a) in Section V.1, followed by our extension to multiple-photon processing (using the conversion from second to first quantization) in Section V.2.
Finally, we summarize our results and discuss potential future works in Section VI.
In the appendices, we propose another symmetrization algorithm, which uses a slightly deeper (poly-logarithmic) circuit but achieves a higher output fidelity. In particular, we present a new symmetrization algorithm for strictly increasing input lists in Appendix A, which is used as a subroutine in the general symmetrization algorithm described in Appendix B.
@@ I.1 Problem setup
We design an $O(\log n)$-depth quantum algorithm for symmetrizing general increasing lists that may include repeated elements, which we call non-strictly increasing lists (NSILs). Recall that this symmetrization problem has been solved for strictly increasing lists (SILs), which do not have any repetition. We formally define such lists as follows.
@@ I.2 Our results
We provide the first two algorithms for symmetrizing NSILs. These algorithms share a common framework, using SIL symmetrization algorithms as subroutines.
The first result is an algorithm with $O(\log n)$ depth and arbitrarily small infidelity. The Berry et al. (2018) algorithm is modified to work as a subroutine of this algorithm. Note that $m$ is the greatest possible value in the NSIL.
Result 1 ($O(\log n)$-depth NSIL symmetrization, Sections III.3 and III.4).
For all $a\geq 2$, there exists an $O(\log n)$-depth NSIL symmetrization algorithm using $O(an\log n+\log m)$ ancilla qubits that produces symmetrized states with fidelity $1-O(n^{-a+2})$.
It is worth noting that, if the input state is simply |$\ket{l}$, rather than a superposition of different |$\ket{l}$s, then the symmetrized state can be prepared with fidelity $1$.
Our second algorithm has slightly greater depth ($O(\log^{3}n)$) but features zero infidelity. It uses a novel SIL symmetrization algorithm as a subroutine, which is based on quantization of a classical parallel algorithm for generating random permutations Alonso and Schott (1996). We postpone the full description of this result to the appendix.
Result 2 (Exact $O(\log^{3}n)$-depth NSIL symmetrization, Appendices A and B).
There exists an $O(\log^{3}n)$-depth NSIL symmetrization algorithm using $O(n\log n)$ ancilla qubits that produces symmetrized states with fidelity $1$.
We also propose an $O(\log^{2}n)$-depth algorithm to convert a list of occupation numbers to an $\mathrm{NSIL}$ of the occupied modes. Combining this algorithm and the symmetrization algorithm, we introduce the first polylogarithmic-depth quantum algorithm to transform a second-quantized state to a first-quantized state. Since this algorithm is a unitary, one can also transform from a first-quantized state to a second-quantized state by running it backward.
Result 3 ($O(\log^{2}n)$-depth conversion between second quantization and first quantization, Section IV).
For all $a\geq 2$, there exists an $O(\log^{2}n)$-depth algorithm using $O(an\log n+\log m)$ ancilla qubits that transforms a state of $n$ particles in $m$ modes from a second-quantized representation to a first-quantized representation, or from a first-quantized representation to a second-quantized representation, with fidelity $1-O(n^{-a+2})$.
@@ I.3 Applications
We review the history of initial state preparation for first-quantized bosonic systems. Abrams and Lloyd (1997) mainly considered Fermi systems and described a $\operatorname{poly}(n)$-depth symmetrization algorithm for SILs, which is improved by Ward et al. (2009). Although both papers claimed to be applicable for bosons, we explain why this is not the case in Sections III.1 and III.2. More recently, Berry et al. (2018) proposed the first $O(\log n)$-depth SIL symmetrization algorithm, as discussed in Sections II.2 and III.2. Several papers proposed using these three algorithms to simulate bosons, without noticing the subtlety caused by repeated items. This includes the recent work of Su et al. (2021) which analyzed resource requirements for quantum chemistry simulation in first quantization and claimed that their algorithms also apply to bosons. Furthermore, Chan et al. (2023); Georges et al. (2024) discussed quantum chemistry simulation and suggested that initial state preparation for bosons can be done in a similar way as for fermions.
There are also several other recent theoretical developments in first-quantized simulation of bosons that do not discuss initial state preparation, including Berry et al. (2023), which discussed simulation of materials using the strategy in Su et al. (2021); and Mukhopadhyay et al. (2024), which studies simulation of Pauli-Fierz Hamiltonians. Similarly, Tong et al. (2022) discussed simulation of gauge theories and bosonic systems, but the authors did not consider the computational cost of the initial state preparation.
As the most straightforward application of Results 1 and 2, the NSIL symmetrization algorithms enable efficient initial state preparation for first-quantized bosonic simulation. Furthermore, as mentioned in Ward et al. (2009), one can also prepare a first-quantized initial state by preparing a second-quantized initial state and transforming it to its first-quantized counterpart. Our Result 3 implements this transformation in poly-logarithmic depth. This idea is particularly useful if the second-quantized state is easier to prepare than the superposition of NSILs.
Although there are currently fewer proposals for first-quantized simulation of bosons than of fermions, our algorithm unlocks future possibilities in this direction if first quantization is shown to be more efficient in, for example, simulation of Bose-Hubbard models Kuwahara et al. (2024) or nuclear effective theory Watson et al. (2023).
@@
Bärtschi and Eidenbenz (2019) $O(n)$ $1$ 00 LNN
Bärtschi and Eidenbenz (2022) $O(n)$ $1$ 00 LNN and all-to-all
Buhrman et al. (2023) $O(\log n)$ $1$ $\Omega(n^{2})$ LNN
Piroli et al. (2024) $O(1)$ $O(\sqrt{n})$ $O(n\log n)$ LNN
Piroli et al. (2024) $O(\log n)$ $O(\sqrt{n})$ $O(n)$ LNN
Result 1 $O(\log n)$ $O(1)$ $O(n\log n)$ All-to-all
Result 2 $O(\log^{3}n)$ $1$ $O(n\log n)$ All-to-all
Table 1: Comparison of Dicke state preparation algorithms for Hamming weight $k=O(n)$. Note that the algorithm of Piroli et al. (2024) must be repeated $O(\sqrt{n})$ times, which implies either a larger number of ancilla qubits ($O(n^{3/2}\cdot\operatorname{poly}(\log n))$) or polynomial total circuit depth.
@@
The Dicke state |$\ket{D_{n}^{k}}$ Dicke (1954) is the equal superposition of all $n$-qubit computational-basis states with Hamming weight $k$, i.e.,
@@
Dicke states play an important role in various areas of quantum information science, including quantum networks Prevedel et al. (2009), quantum metrology Tóth (2012), quantum error correction Ouyang (2014), and the quantum approximate optimization algorithm Hadfield et al. (2019). Therefore, the quantum information community has sought circuits for preparing Dicke states with lower depth, gate complexity, and number of ancilla qubits. The most challenging Dicke states to prepare are those with $k$ close to $n/2$. Linear-depth circuits for preparing these Dicke states were proposed in Refs. Bärtschi and Eidenbenz (2019, 2022), but the first $\operatorname{poly}(\log n)$-depth circuits were proposed by Buhrman et al. (2023). These circuits use a model of “local alternating quantum classical computations” in which the quantum operations performed are determined by the measurement outcomes of the previous step. However, the number of ancilla qubits required in that algorithm is huge ($\Omega(n^{2})$) for large $n$. Piroli et al. (2024) significantly reduce the number of ancilla qubits and circuit depth to $O(n\log n)$ and $O(1)$, respectively, or $O(n)$ and $O(\log n)$, respectively. However, the circuit in Ref. Piroli et al. (2024) needs to be repeated $O(\sqrt{n})$ times in total, which implies either a larger number of ancilla qubits $O(n^{3/2}\cdot\operatorname{poly}(\log n))$ or polynomial total circuit depth.
Our low-depth NSIL symmetrization algorithm gives a straightforward strategy for preparing arbitrary Dicke states with $O(\log n)$ depth and $O(n\log n)$ ancilla qubits, with the cost of increased circuit connectivity. To prepare |$\ket{D_{n}^{k}}$, it suffices to symmetrize
@@
using our algorithm in Section III.3. Since the initial state is only one NSIL, it suffices to use the single-input symmetrization algorithm.
Superpositions of different Dicke states (with different $k$ values) are also useful in quantum metrology. For example, Lin et al. (2024) show that metrological entanglement advantage can be achieved with such states. We can prepare such states in $O(\log n)$ depth by first initializing the state as a superposition of different |$\ket{\psi_{n,k}}$ states and then applying the full NSIL symmetrization algorithm described in Section III.4. While some Dicke state preparation circuits can straightforwardly prepare superpositions of Dicke statesBärtschi and Eidenbenz (2019, 2022), this is not the case for previous polylogarithmic-depth circuits Buhrman et al. (2023); Piroli et al. (2024), which may need to be modified and might incur extra overhead.
@@
As a promising candidate to achieve microarcsecond resolution in astronomical observation, long-baseline telescopes based on quantum repeaters were first proposed in Gottesman et al. (2012) and have sparked increasing research interest in both the astronomy and quantum information communities. In follow-up work, Khabiboulline et al. (2019a) design a quantum telescope array and employ the quantum Fourier transform (QFT) to improve the signal-to-noise ratio. However, their method can only handle the case where the number of photons per mode is restricted to be much less than $1$ (i.e., the object must be extremely dim). Using our NSIL symmetrization algorithm, we extend the QFT-based scheme to cases where some modes have more than 1 photon. Since symmetrization is an intermediate step of interferometric imaging, we postpone the detailed description of this application to Section V.

### v2

@@ Abstract
Quantum symmetrization is the task of transforming a non-strictly increasing list of $n$ integers into an equal superposition of all permutations of the list (or more generally, performing this operation coherently on a superposition of such lists). This task plays a key role in initial state preparation for first-quantized simulations. Motivated by an application to fermionic systems, various algorithms have been proposed to solve a weaker version of symmetrization in which the input list is strictly increasing, but the general symmetrization problem with repetitions in the input list has not been well studied. We present the first efficient quantum algorithms for the general symmetrization problem. If $m$ is the greatest possible value of the input list, our first algorithm symmetrizes any single classical input list using $\tilde{O}(\log n)$ depth and $O(n\log n+\log m)$ ancilla qubits, and our second algorithm symmetrizes an arbitrary superposition of input lists using $\tilde{O}(\log^{3}n)$ depth and $O(n\log n)$ ancilla qubits. Our algorithms enable efficient simulation of bosonic quantum systems in first quantization and can prepare (superpositions of) Dicke states of any Hamming weight in $\tilde{O}(\log n)$ depth (respectively, $\tilde{O}(\log^{3}n)$ depth) using $O(n\log n)$ ancilla qubits. We also propose an $\tilde{O}(\log^{3}n)$-depth quantum algorithm to transform second-quantized states to first-quantized states. Using this algorithm, QFT-based quantum telescope arrays can image brighter photon sources, extending quantum interferometric imaging systems to a new regime.
@@ I Introduction
Symmetry is a central concept in physics. In particular, symmetry plays a key role in quantum simulation [1], where a faithful simulation of the many-body dynamics of indistinguishable particles should respect the symmetry of the simulated system. For fermions, due to the Pauli exclusion principle, no two particles can occupy the same mode, and the wave function must be antisymmetrized with respect to permutations of particles, leading to Fermi-Dirac statistics. For bosons, multiple particles can be in the same state, and the wave function must be symmetrized with respect to permutations, leading to Bose-Einstein statistics. A general quantum simulation procedure contains not only Hamiltonian evolution and final output, but also initial state preparation, which is often a challenging step, especially when the initial wave function has some specific structure or symmetry.
@@
In this paper, we study a simple but subtle problem in quantum computing: transforming a list $l$ of non-strictly increasing integers stored in a quantum register into the equal superposition of all permutations of $l$. We call this “symmetrization,” since the output state is symmetric with respect to any permutation of $l$. Symmetrization plays a fundamental role in quantum simulation based on first quantization, an encoding of identical particles as a list of particle locations (whereas second quantization represents the state as a list of occupation numbers of particle locations). Eigenstates in first quantization must be symmetric or antisymmetric if the system is bosonic or fermionic, respectively. Due to significant interest in quantum chemistry (which involves interactions between fermions), a weaker version of (anti)symmetrization where the list is strictly increasing has been extensively studied [2, 3, 4]. In particular, work of Berry et al. [4] develops a logarithmic-depth algorithm. However, first-quantized simulation of bosons must allow multiple particles to be in the same mode, so the input list may contain repeated integers. Although several previous papers have claimed to handle this more general case [2, 3, 5, 6, 7, 8], close examination reveals a subtlety that has not been properly addressed in the literature. The only solution for the general symmetrization problem that we are aware of is a result of Nepomechie and Raveh [9]. However, this algorithm requires prohibitively deep circuits: the depth is $O(n^{m})$, where $n$ is the number of integers and $m$ is the greatest of these integers.
@@
As the main results of this paper, we propose new logarithmic-depth quantum algorithms for symmetrization and a polylogarithmic-depth quantum algorithm for transforming from second to first quantization. We also present three applications of these algorithms. The first application, mentioned above, is initial state preparation in first-quantized simulation of bosons, an unrecognized unsolved problem before our work. The second application is preparing a specific class of symmetric states useful in a variety of quantum information processing tasks, Dicke states, where our algorithm is the first  [^1] We are aware of concurrent work that also gives an algorithm for preparing Dicke states with polylogarithmic depth but uses only polylogarithmically many ancillas [37]. The approaches are very different: while we use sorting networks, the other approach applies a simple sequence of collective rotations and parity measurements. Our approach solves a more general symmetrization problem, while the other approach is simpler, likely performs better in practice, and can perform better when the Hamming weight is lower. to achieve logarithmic total circuit depth using a reasonable number of ancilla qubits. The third application is in optics, where we use our algorithm to improve strategies for long-baseline quantum telescope arrays [11, 12] to process multiple photons at a time, yielding a quantum interferometric imaging system that can take images in more general scenarios than previous designs.
@@
We present background for our results in Section II. More specifically, in Section II.1, we introduce sorting networks and their quantization as the building blocks of symmetrization. Then in Section II.2, we describe the algorithm of Berry et al. [4], which is used as a subroutine of one of our symmetrization algorithms. In Section II.3, we generalize sorting to different comparison rules and define reversible sorting networks. In Section II.4, we present the quantization of a standard prefix sum computation procedure.
In Section III, we discuss the general symmetrization problem with repetitions. We first explain in Section III.1 why previous methods [2, 3, 4] are not applicable to the more general symmetrization problem by discussing the subgroup structure of permutations when considering repetitions in the permuted list, and then extend quantum sorting networks to work with repetitions in Section III.2. Using these theoretical tools, we present our new algorithm for symmetrizing a single input list in Section III.3.
We present our symmetrization algorithm for an arbitrary superposition of input lists in Section IV. We present a novel SIL symmetrization procedure in Section IV.1 based on quantization of a classical parallel algorithm from Alonso and Schott [13], which is subsequently used as a subroutine in Section IV.2 to design the NSIL symmetrization algorithm for superposed inputs.
In Section V, we propose a polylogarithmic-depth converter from a second quantized representation to a sorted list of modes occupied by the particles. This binary encoder, together with the symmetrization algorithm, gives the first polylogarithmic-depth conversion from second quantization to first quantization.
As a topic of independent interest, we demonstrate our design of a quantum interferometric imaging system in Section VI, including a pedagogical introduction of quantum telescope arrays that process one photon a time [11, 12] in Section VI.1, followed by our extension to multiple-photon processing (using the conversion from second to first quantization) in Section VI.2.
Finally, we summarize our results and discuss potential future works in Section VII.
@@ I.1 Problem setup
We design an $\tilde{O}(\log n)$-depth quantum algorithm for symmetrizing general increasing lists that may include repeated elements, which we call non-strictly increasing lists (NSILs). Here we define $\tilde{O}(f(n))\coloneqq O(f(n)\log(f(n)))$. Note that we will use this notation with $f(n)$ scaling like $\log n$, in which case the $\tilde{O}$ notation neglects a factor of $\log\log n$.
Recall that this symmetrization problem has been solved for strictly increasing lists (SILs), which do not have any repetition. We formally define such lists as follows.
@@ I.2 Our results
We provide two NSIL symmetrization algorithms, one for single NSIL inputs and another for superposed NSIL inputs. Both algorithms use SIL symmetrization algorithms as subroutines.
The first result uses a modified version of the Berry et al. [4] algorithm and is designed for a single NSIL input. It takes $\tilde{O}(\log n)$ depth.
Result 1 ($\tilde{O}(\log n)$-depth single-input NSIL symmetrization, Section III.3).
For all $a\geq 2$, there exists an $\tilde{O}(\log n)$-depth algorithm using $O(an\log n+\log m)$ ancilla qubits that produces |$\ket{\psi_{\mathrm{out}}(l)}$ for any $l\in\mathrm{NSIL}_{n}$ with success probability $1-O(n^{-a+2})$.
Our second algorithm has slightly greater depth ($\tilde{O}(\log^{3}n)$) but works for arbitrary superpositions of inputs $\sum_{l\in\mathrm{NSIL}_{n}}\alpha_{l}\ket{l}$ and always succeeds. It uses a novel SIL symmetrization algorithm as a subroutine, which is based on quantization of a classical parallel algorithm for generating random permutations [13].
Result 2 ($\tilde{O}(\log^{3}n)$-depth superposed-input NSIL symmetrization, Section IV).
There exists an $\tilde{O}(\log^{3}n)$-depth NSIL symmetrization algorithm using $O(n\log n)$ ancilla qubits that produces $\sum_{l\in\mathrm{NSIL}_{n}}\alpha_{l}\ket{\psi_{\mathrm{out}}(l)}$ for any input state $\sum_{l\in\mathrm{NSIL}_{n}}\ket{\psi_{\mathrm{out}}(l)}$ with success probability $1$.
We also propose an $\tilde{O}(\log^{2}n)$-depth algorithm to convert a list of occupation numbers to an $\mathrm{NSIL}$ of the occupied modes. Combining this algorithm and Result 2, we introduce the first polylogarithmic-depth quantum algorithm to transform a second-quantized state to a first-quantized state. Since this algorithm is a unitary, one can also transform from a first-quantized state to a second-quantized state by running it backward.
Result 3 ($\tilde{O}(\log^{3}n)$-depth conversion between second quantization and first quantization, Section V).
There exists an $\tilde{O}(\log^{3}n)$-depth algorithm using $O(n\log n)$ ancilla qubits that transforms a state of $n$ particles in $m$ modes from a second-quantized representation to a first-quantized representation, or from a first-quantized representation to a second-quantized representation.
@@ I.3 Applications
We review the history of initial state preparation for first-quantized bosonic systems. Abrams and Lloyd [2] mainly considered Fermi systems and described a $\operatorname{poly}(n)$-depth symmetrization algorithm for SILs, which is improved by Ward et al. [3]. Although both papers claimed to be applicable for bosons, we explain why this is not the case in Sections III.1 and III.2. More recently, Berry et al. [4] proposed the first $\tilde{O}(\log n)$-depth SIL symmetrization algorithm, as discussed in Sections II.2 and III.2. Several papers proposed using these three algorithms to simulate bosons, without noticing the subtlety caused by repeated items. This includes the recent work of Su et al. [5] which analyzed resource requirements for quantum chemistry simulation in first quantization and claimed that their algorithms also apply to bosons. Furthermore, Chan et al. [6], Georges et al. [8] discussed quantum chemistry simulation and suggested that initial state preparation for bosons can be done in a similar way as for fermions.
There are also several other recent theoretical developments in first-quantized simulation of bosons that do not discuss initial state preparation, including Berry et al. [14], which discussed simulation of materials using the strategy in Su et al. [5]; and Mukhopadhyay et al. [15], which studies simulation of Pauli-Fierz Hamiltonians. Similarly, Tong et al. [16] discussed simulation of gauge theories and bosonic systems, but the authors did not consider the computational cost of the initial state preparation.
As the most straightforward application of Results 1 and 2, the NSIL symmetrization algorithms enable efficient initial state preparation for first-quantized bosonic simulation. Furthermore, as mentioned in [3], one can also prepare a first-quantized initial state by preparing a second-quantized initial state and transforming it to its first-quantized counterpart. Our Result 3 implements this transformation in poly-logarithmic depth. This idea is particularly useful if the second-quantized state is easier to prepare than the superposition of NSILs.
Although there are currently fewer proposals for first-quantized simulation of bosons than of fermions, our algorithm unlocks future possibilities in this direction if first quantization is shown to be more efficient in, for example, simulation of Bose-Hubbard models [17] or nuclear effective theory [18].
@@
Bärtschi and Eidenbenz [19] $O(n)$ $1$ 00 LNN
Bärtschi and Eidenbenz [20] $O(n)$ $1$ 00 LNN and all-to-all
Buhrman et al. [21] $O(\log n)$ $1$ $\Omega(n^{2})$ LNN
Piroli et al. [22] $O(1)$ $O(\sqrt{n})$ $O(n\log n)$ LNN
Piroli et al. [22] $O(\log n)$ $O(\sqrt{n})$ $O(n)$ LNN
Result 1 $\tilde{O}(\log n)$ $O(1)$ $O(n\log n)$ All-to-all
Result 2 $\tilde{O}(\log^{3}n)$ $1$ $O(n\log n)$ All-to-all
Table 1: Comparison of Dicke state preparation algorithms for Hamming weight $k=O(n)$. Note that the algorithm of Piroli et al. [22] must be repeated $O(\sqrt{n})$ times, which implies either a larger number of ancilla qubits ($O(n^{3/2}\cdot\operatorname{poly}(\log n))$) or polynomial total circuit depth.
@@
The Dicke state |$\ket{D_{n}^{k}}$ [23] is the equal superposition of all $n$-qubit computational-basis states with Hamming weight $k$, i.e.,
@@
Dicke states play an important role in various areas of quantum information science, including quantum networks [24], quantum metrology [25], quantum error correction [26], and the quantum approximate optimization algorithm [27]. Therefore, the quantum information community has sought circuits for preparing Dicke states with lower depth, gate complexity, and number of ancilla qubits. The most challenging Dicke states to prepare are those with $k$ close to $n/2$. Linear-depth circuits for preparing these Dicke states were proposed in Refs. [19, 20], but the first $\operatorname{poly}(\log n)$-depth circuits were proposed by Buhrman et al. [21]. These circuits use a model of “local alternating quantum classical computations” in which the quantum operations performed are determined by the measurement outcomes of the previous step. However, the number of ancilla qubits required in that algorithm is huge ($\Omega(n^{2})$) for large $n$. Piroli et al. [22] significantly reduce the number of ancilla qubits and circuit depth to $O(n\log n)$ and $O(1)$, respectively, or $O(n)$ and $O(\log n)$, respectively. However, the circuit in Ref. [22] needs to be repeated $O(\sqrt{n})$ times in total, which implies either a larger number of ancilla qubits $O(n^{3/2}\cdot\operatorname{poly}(\log n))$ or polynomial total circuit depth.
Our low-depth NSIL symmetrization algorithm gives a straightforward strategy for preparing arbitrary Dicke states with $\tilde{O}(\log n)$ depth and $O(n\log n)$ ancilla qubits, with the cost of increased circuit connectivity. To prepare |$\ket{D_{n}^{k}}$, it suffices to symmetrize
@@
using our algorithm in Section III.3, since the initial state is only one NSIL.
Superpositions of different Dicke states (with different $k$ values) are also useful in quantum metrology. For example, Lin et al. [28] show that metrological entanglement advantage can be achieved with such states. We can prepare such states in $\tilde{O}(\log^{3}n)$ depth by first initializing the state as a superposition of different |$\ket{\psi_{n,k}}$ states and then applying the full NSIL symmetrization algorithm described in Section IV. While some Dicke state preparation circuits can straightforwardly prepare superpositions of Dicke states[19, 20], this is not the case for previous polylogarithmic-depth circuits [21, 22], which may need to be modified and might incur extra overhead.
@@
As a promising candidate to achieve microarcsecond resolution in astronomical observation, long-baseline telescopes based on quantum repeaters were first proposed in Gottesman et al. [11] and have sparked increasing research interest in both the astronomy and quantum information communities. In follow-up work, Khabiboulline et al. [12] design a quantum telescope array and employ the quantum Fourier transform (QFT) to improve the signal-to-noise ratio. However, their method can only handle the case where the number of photons per mode is restricted to be much less than $1$ (i.e., the object must be extremely dim). Using our NSIL symmetrization algorithm, we extend the QFT-based scheme to cases where some modes have more than 1 photon. Since symmetrization is an intermediate step of interferometric imaging, we postpone the detailed description of this application to Section VI.